OpenStudy (anonymous):

The function s(t)=t^2-3t-4 describes how the speed of a ball (S, measured in feet per second) changes with respect to time (t, measured in seconds) as it rolls down a hill. How many seconds does it take for the ball to stop?

OpenStudy (jim766):

you need to factor it, do you know how?

OpenStudy (campbell_st):

well you need to let s(t) = 0.... no speed and then solve for t so \[t^2 -3t -4 = 0\] only consider the positive solution for t... as time can't be negative. and the quadratic can be factored.

OpenStudy (anonymous):

-4 and 1

OpenStudy (campbell_st):

not quite \[t^2 -3t -4 = 0... factors ..\to... (t -4)(t + 1)=0\] so you need to now solve t - 4 = 0 and t + 1 = 0

OpenStudy (anonymous):

That's was I said. I just gave you the factors

OpenStudy (anonymous):

It took the ball 4 seconds to stop then.

OpenStudy (campbell_st):

that seems to be the answer if you substitute it into your initial equation you should find s(t) = 0

OpenStudy (anonymous):

Yes. Thanks a bunch!